OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 4 sin πt + 5 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) Find the average velocity during each time period. 1,2 1,1.1 1,1.01 1,1.001???

OpenStudy (anonymous):

Velocity is the derivative of displacement with respect to time. Your first step therefore is to differentiate s with respect to t!

OpenStudy (anonymous):

so derive the equation??

OpenStudy (anonymous):

Yeps. Give it a shot.

OpenStudy (anonymous):

omg o.O ive never derived sin or cos before

OpenStudy (anonymous):

pi(4cospit)-(5sinpit)?

OpenStudy (anonymous):

That's almost right, good job for your first try at sin and cos :) In fact, I suspect it's a typo, or just an oversight that you wrote \(5sin(\pi t\)) instead of \(5\pi sin(\pi t)\)

OpenStudy (anonymous):

The rest is correct, including your signs. Now I'm not entirely sure about the rest of the question. What I can tell you is that sin and cos repeat themselves from 0-2\(\pi\) (one period). Since the question asks about time periods, I'm guessing that you need to find the velocity at 0, and at 2\(\pi\). These would be at t=0, and t=2 respectively. The average would then probably be the sum divided by two (though there may be some calculus that I'm missing).

OpenStudy (anonymous):

mmm okay i think i get it. thank you for helping me! could you maybe help me with one more?

OpenStudy (anonymous):

i have the same assignment, and we need to plug in the numbers and guess, not differentiate because we havent done that yet in class... but I still cant do it...

OpenStudy (anonymous):

did you get it right christaljade?